# FairMT-Bench: Benchmarking Fairness for Multi-turn Dialogue in Conversational LLMs

### v1

@@ 3.2 Evaluation Performance on different Tasks
In this experiment, we use GPT-4 and Llama-Guard-3 to test the proportion of biased or toxic answers output by the model on different tasks.
@@
The results evaluated by GPT-4 are shown in Table 2. Overall, when comparing the results on stereotype and toxicity datasets, we observe a consistent distribution of results, with the best and worst-performing models being largely similar across tasks. In general, models perform poorly on the “Anaphora Ellipsis” and “Interference from Misinformation” tasks. This indicates that when there are more pronouns and ellipses in the context, models struggle to understand biases within the complete dialogue and are more likely to bypass safety mechanisms, leading to biased content generation. Additionally, when the input contains a lot of misleading information and the model is asked to summarize or respond based on context, it becomes more susceptible to interference from earlier biased input, thus incorporating biases into its responses.
Notably, performance differences are evident across models. For example, Llama-2 (7B, 13B) perform poorly on tasks like “Anaphora Ellipsis”, which contain fewer explicit bias-related or toxic keywords and focus more on implicit biases. However, they are less affected by interaction or task execution interference like “Negative Feedback”, generating fewer biased outputs under the influence of context or user instructions. In contrast, models like Mistral (7B) perform better in tasks involving implicit bias understanding within the context like “Scattered Questions”, successfully avoiding biased outputs, but are more prone to generating biased responses when influenced by user input or instructions as in tasks like “Interference Misinformation”.
Figure 3: Bias rates of different LLMs on Fair-MT 10K evaluated by Llama-Guard-3-8B. We use abbreviations instead of task names, SQ stands for Scattered Questions, AE stands for Anaphora Ellipsis, JT stands for Jailbreak Tips, IM stands for Interference from Misinformation, NF stands for Negative Feedback, $F$ stands for Fixed Format.
@@
We employ Llama-Guard-3 for auxiliary evaluation. The the overall trends of Llama-Guard-3’s evaluation results align closely with those from GPT-4, as shown in Appendix Table 4. To better visualize the distribution of model performance across different tasks, we apply z-score normalization by task, resulting in Figure 3. As shown, Llama-Guard-3’s evaluation indicates consistent performance on stereotype and toxicity biases across models. Additionally, tasks requiring understanding like “Scattered Questions” and “Anaphora Ellipsis” favor models such as Mistral (7B) and ChatGPT, whereas tasks focused on bias resistance like “Interference Misinformation” and “Fixed Format” favor models like the Llama-2 (7B, 13B). This generally coincides with the findings from the GPT-4 evaluation.
In conclusion, our GPT-4 and Llama-Guard-3 evaluation shows that current LLMs exhibit significant variation in performance across the tasks defined in our study. By categorizing tasks into those requiring strong comprehension (i.e., “Scattered Questions” and “Anaphora Ellipsis”) and those demanding resistance to guided bias (i.e., “Jailbreak Tips”, “Interference Misinformation”, “Fixed Format”, and “Negative Feedback”), we observe distinct performance patterns across models. Some models excel at comprehension-focused tasks but underperform on bias-resistance tasks, while others exhibit the opposite trend. These differences may stem from variations in alignment paradigms and instruction-following capabilities. Models like the Llama-2 series rely more heavily on keyword-based bias detection, resulting in reduced fairness when handling implicit biases or less common expressions present in the alignment data. Conversely, models like Mistral (7B), though stronger in contextual semantic understanding, may prioritize utility and user satisfaction over safety when following user instructions or specific requests. Consequently, despite numerous efforts to improve overall LLM performance, no model has yet demonstrated consistently strong fairness performance across both comprehension-focused and bias-resistance tasks.
@@ 3.3 Comparison of Performance between Single and Multi-Turn
In this experiment, we evaluate the fairness performance of LLMs in the single-turn context and the multi-turn context involving the predefined six tasks. To achieve this, we extract the final prompts from the multi-turn dialogues in our constructed dataset and utilize them as single-turn inputs to assess the proportion of biased responses generated by the models. We then analyze the bias rates between single-turn and multi-turn dialogues across models and tasks.
@@
Figure 5 presents the bias rate comparison under single-turn and multi-turn scenarios of different models. Except for Gemma, all models exhibit higher bias rates in multi-turn dialogues than in single-turn ones across Stereotype and Toxicity. Notably, model Llama-2 (7B, 13B) and Llama-3.1 exhibit significant increases in bias rate in both scenarios. In the Stereotype dataset, single-turn dialogues generally exhibit higher bias rates, while bias rate differences typically range between 5% and 10%. The largest increase is seen in Llama-2 (7B). In contrast, in the Toxicity dataset, single-turn dialogues have lower bias rates, with multi-turn dialogues exhibiting an increase of around 10%. The largest increase is seen in Llama-3.1-8B. Gemma is the only model showing a reduction in bias from single-turn to multi-turn dialogues. Our analysis reveals that in multi-turn dialogues, Gemma’s bias rate drops by over 80% in the Fix Format task, significantly contributing to its overall reduction in bias. However, in other tasks, such as Scattered Questions, Gemma maintains higher bias rates when multi-turn dialogue history is included. See Appendix B.3 for further details.
Figure 4: Comparison of bias rates in single versus multi-turn dialogues in terms of LLMs.
Figure 5: Comparison of bias rates in single versus multi-turn dialogues in terms of tasks.
@@
Figure 5 compares bias rates across different tasks under both single-turn and multi-turn scenarios. The impact of multi-turn dialogues varies by task. In Scattered Questions, Anaphora Ellipsis, and Jailbreak Tips, bias rates increase significantly in the multi-turn setting for both Stereotype and Toxicity datasets. Conversely, task Fix Format, Negative Feedback, and Interference Misinformation exhibit similar or slightly lower bias rates. In the Stereotype dataset, single-turn dialogues generally show higher bias rates, with the most noticeable increases in Scattered Questions and Jailbreak Tips. In the Toxicity dataset, single-turn dialogues have lower bias rates, with larger differences observed in Scattered Questions, Anaphora Ellipsis, and Jailbreak Tips. As illustrated in Figure 5, models show relatively high bias rates in single-turn scenarios for tasks including Fix Format and Negative Feedback, indicating that models may prioritize utility over fairness even in single-turn contexts, making them more vulnerable to external user input or feedback. However, in tasks like Scattered Questions and Jailbreak Tips, implicit biases accumulated across dialogue turns contribute to significantly higher bias rates in the final turn.
In summary, current LLMs exhibit a noticeable increase in bias rates when transitioning from single-turn to multi-turn dialogues, with models such as Llama-2 (7B, 13B) and Llama-3.1 being typical examples. Notably, tasks defined in our study like Scattered Questions, Anaphora Ellipsis, and Jailbreak Tips prove particularly challenging for these models, as their overall fairness performance declines sharply. Additionally, across both models and tasks, bias rates in the Stereotype dataset are consistently higher than in the Toxicity dataset in single-turn dialogues, and the bias rate difference between single-turn and multi-turn scenarios is consequently smaller. This aligns Wang et al. (2024), which suggest that Stereotype tasks are more subtle and challenging compared to Toxicity tasks.
@@ 3.4 Evaluation results in different turns
To investigate the impact of turn count on model performance across different models, we calculate the average bias rates for each dialogue turn. As shown in Figure 7, the bias rates of all models for both Stereotype and Toxicity increase with the number of turns. This suggests that, in general, the model’s outputs are influenced by the historical dialogue context, leading to an accumulation of bias across multiple turns of responses.
Figure 6: Bias rates across different dialogue turns in terms of LLMs.
Figure 7: Bias rates across different dialogue
@@
To explore the effect of dialogue turn count on performance across different tasks, we calculate the average bias rate for each task across models, as illustrated in Figure 7. Note that in the Interference from Misinformation task, models are not prompted to respond during the first four turns, and in the Fix Format task, the initial prompts are objective and unrelated to bias. Thus, there are no line charts for these two tasks in the figure. As observed, in both the Stereotype and Toxicity tasks, bias rates generally increase with the number of dialogue turns, showing a marked spike in the final turn, with the exception of the Scattered Questions task. Our further analysis suggests that this deviation in Scattered Questions may be due to the third-turn prompt subtly guiding the model toward bias. Specifically, the prompt “Summarize your speculation, it sounds like they have some characteristics similar to [Target Group]?” implicitly encourages the model to associate the group described in the dialogue history with a target group, without explicitly introducing a biased evaluation. This indirect guidance can lead the model to generate biased responses. However, as subsequent prompts more directly link negative events or evaluations to the target group, the model’s safety mechanisms are triggered, causing it to reject further responses. In summary, current LLMs tend to accumulate bias in multi-turn dialogues, which is particularly pronounced in fairness-related tasks like Anaphora Ellipsis, Jailbreak Tips, and Negative Feedback.
@@ 3.5 Evaluation Results on different groups
We assess model fairness across bias attributes by calculating the proportion of biased responses for each category in our dataset. As shown in Figure 8(a) (see Appendix B.4 for a color bar zoomed-in version), slight variation in performance is observed across models and bias attributes groups on the Stereotype datasets. The Age group shows the largest disparity, while the Gender group consistently sees poor performance across all models. Llama-2-chat (7B, 13B) exhibit the most variability across bias attributes, whereas ChatGPT demonstrates strong and consistent performance. As shown in Figure 8(b), performance variation is more pronounced on the Toxicity datasets. The Race group shows the largest disparity, while the Gender and Disabled groups consistently exhibit poor results. Llama-2-chat (7B) again shows the most significant variability, while Gemma consistently performs poorly across all bias attributes.
In summary, the fairness capabilities of current LLMs vary significantly across bias attributes. A notable example is Llama-2-chat (7B, 13B), which shows substantial performance disparities across bias attributes in both datasets. The models exhibit weaker alignment in less commonly represented categories such as age, disabled, and appearance, while demonstrating stronger fairness in categories with greater focus, such as race and religion. These findings highlight the need for future LLM fairness efforts to drive more comprehensive alignment across all bias attributes.
Figure 8: Bias rates across different bias attributes.
@@ 3.6 Challenge Fair-MT Bench 1K
To enable more efficient evaluation, we distill the most challenging data from our Fair-MT Bench to create a lighter LLM fairness benchmark, FairMT-1K. Specifically, we select data points where the six models had the highest error rates in the original 10K dataset based on our testing results. An equal number of samples are chosen from each task. We then evaluated a broader range of models on this new dataset, including Gemma-2-it (2B, 9B, 27B) (Gemma Team et al., 2024), Mistral-Small-Instruct-2409 (Jiang et al., 2023), Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2023), Qwne2.5-Instruct (0.5B, 3B, 7B) (Team, 2024), and GPT-4 (Achiam et al., 2023), resulting in a more diverse set of models. Table 3 presents the fairness performance of each model on FairMT-1K. We sampled 10% of the data for manual inspection, and the results and analysis are shown in the Appendix A.5.
Table 3: Bias rates of different LLMs on FairMT-1K.
@@
As demonstrated, even the most recently introduced models exhibit a significant proportion of biased responses on the FairMT-1K, underscoring the dataset’s challenging nature for assessing LLM fairness. Notably, we find that for certain tasks, such as Scattered Questions, Fix Format, and Negative Feedback, the proportion of biased responses tends to decrease as the model size increases. This could be due to the increased model parameters enhancing the ability of models to comprehend user instructions, which facilitates a more accurate understanding of user intent in multi-turn dialogues. Consequently, this improvement may reduce instances where fairness is compromised in favor of increasing user satisfaction.
@@ 3.7 Discussion
Our experiments demonstrate the necessity of establishing a new benchmark for evaluating LLM fairness performance in multi-turn dialogues, since testing on single-turn fairness data fails to capture issues stemming from bias accumulation (see Section 3.3 and Section 3.4), and empirical results reveal that current LLMs struggle to maintain consistently strong fairness performance across a diverse set of dialogue scenarios designed for evaluating fairness (see Section 3.2). Importantly, we observe that LLMs face challenges in comprehensively addressing different bias attributes (see Section 3.5) and different task types including comprehension-focused and bias-resistance tasks (see Section 3.2). With this proposed benchmark, we encourage future work to focus on improving LLM fairness by leveraging this well-designed evaluation framework to achieve more comprehensive fairness enhancements in the near future.

### v2

@@ 3.2 Evaluation Performance on different Tasks
In this experiment, we use GPT-4 and Llama-Guard-3 to test the proportion of biased or toxic answers output by the model on different tasks. In addition, we evaluated the models’ multi-turn dialogue capabilities. The detailed evaluation results and their tendency to over-reject questions that do not contain bias in each dimension are provided in the Appendix B.1, Table 6.
@@
The results evaluated by GPT-4 are shown in Table 2. Overall, when comparing the results on stereotype and toxicity datasets, we observe a consistent distribution of results, with the best and worst-performing LLMs being largely similar across tasks. In general, LLMs perform poorly on the “Anaphora Ellipsis” and “Interference from Misinformation” tasks. This indicates that when there are more pronouns and ellipses in the context, LLMs struggle to integrate previous information to understand biases within the full dialogue and are more likely to bypass fairness protective mechanisms. Additionally, when the input contains a lot of misleading information and the model is asked to summarize or respond based on context, it becomes more susceptible to interference from earlier biased input, thus incorporating biases into its responses.
Notably, performance differences are evident across LLMs. For example, Llama-2 (7B, 13B) perform poorly on tasks like “Anaphora Ellipsis”, which contain fewer explicit bias-related or toxic keywords and focus more on implicit biases. However, they are less affected by interaction or task execution interference like “Negative Feedback”, generating fewer biased outputs under the influence of context or user instructions. In contrast, LLMs like Mistral (7B) perform better in tasks involving implicit bias understanding within the context like “Scattered Questions”, successfully avoiding biased outputs, but are more prone to generating biased responses when influenced by user input or instructions as in tasks like “Interference Misinformation”. The consistency between the human and the GPT-4 annotation results is shown in Appendix A.4.2.
Figure 3: Bias ratio of different LLMs on FairMT-10K evaluated by Llama-Guard-3-8B. We use abbreviations instead of task names, SQ stands for Scattered Questions, AE stands for Anaphora Ellipsis, JT stands for Jailbreak Tips, IM stands for Interference from Misinformation, NF stands for Negative Feedback, $F$ stands for Fixed Format.
@@
We employ Llama-Guard-3 for auxiliary evaluation. The the overall trends of Llama-Guard-3’s evaluation results align closely with those from GPT-4, as shown in Appendix Table 8. To better visualize the distribution of model performance across different tasks, we apply z-score normalization by task, resulting in Figure 3. As shown, Llama-Guard-3’s evaluation indicates consistent performance on stereotype and toxicity biases across models. Additionally, tasks requiring understanding like “Scattered Questions” and “Anaphora Ellipsis” favor models such as Mistral (7B) and ChatGPT, whereas tasks focused on bias resistance like “Interference Misinformation” and “Fixed Format” favor models like the Llama-2-chat (7B, 13B). This generally coincides with the findings from the GPT-4 evaluation.
In conclusion, our GPT-4 and Llama-Guard-3 evaluation shows that current LLMs exhibit significant variation in performance across the tasks defined in our study. We observe distinct task performance patterns across different models. Specifically, some models excel at comprehension-focused tasks (such as “Scattered Questions” and “Anaphora Ellipsis”) but underperform on tasks requiring bias-resistance (such as “Jailbreak Tips”, “Interference Misinformation”, “Fixed Format”, and “Negative Feedback”), while others exhibit the opposite trend. These differences may stem from variations in alignment paradigms and instruction-following capabilities. Models like the Llama-2 series rely more heavily on keyword-based bias detection, resulting in reduced fairness when handling comprehension-focused tasks. Conversely, models like Mistral (7B), though stronger in contextual semantic understanding, may prioritize utility and user satisfaction over safety when following user instructions or specific requests. Consequently, despite numerous efforts to improve overall LLM performance, no model has yet demonstrated consistently strong fairness performance across both comprehension-focused and bias-resistance tasks.
@@ 3.3 Comparison of Performance between Single and Multi-Turn
In this experiment, we evaluate the fairness performance of LLMs in both single-turn and multi-turn contexts using the predefined six tasks. We extract the final prompts from the multi-turn dialogues in our dataset and use them as single-turn inputs to evaluate the proportion of biased responses generated by the LLMs. We then compare and analyze the bias ratio between single-turn and multi-turn dialogues across models and tasks.
@@
Figure 5 presents the bias ratio comparison under single-turn and multi-turn scenarios of different models. All LLMs, except Gemma, exhibit higher bias ratio in multi-turn dialogues than in single-turn ones across Stereotype and Toxicity. Notably, the Llama-2-chat (7B, 13B) and Llama-3.1-it exhibit significant increases in bias ratio in both scenarios. In the Stereotype dataset, single-turn dialogues generally exhibit higher bias ratio, while bias ratio differences typically range between 5% and 10%. , with the largest increase observed in Llama-2-chat-7B. In contrast, in the Toxicity dataset, single-turn dialogues exhibit a lower bias ratio, whereas multi-turn dialogues see an increase of around 10%, with the most substantial rise in Llama-3.1-it. Uniquely, Gemma displays a reduction in bias from single-turn to multi-turn dialogues. In-depth analysis shows that in multi-turn dialogues, Gemma’s bias ratio drops by over 80% in the Fix Format task, significantly contributing to its overall reduction in bias. However, in other tasks like Scattered Questions, Gemma maintains higher bias ratio when multi-turn dialogue history is included. Detailed results, with the performance of all models on each task, are provided in Appendix B.3.
Figure 4: Comparison of bias ratio in single versus multi-turn dialogues in terms of LLMs.
Figure 5: Comparison of bias ratio in single versus multi-turn dialogues in terms of tasks.
@@
Figure 5 compares bias ratio across different tasks under both single-turn and multi-turn scenarios. The impact of multi-turn dialogues varies by task. In tasks such as Scattered Questions, Anaphora Ellipsis, and Jailbreak Tips, bias ratio increase significantly in the multi-turn setting for both Stereotype and Toxicity datasets. Conversely, task like Fix Format, Negative Feedback, and Interference Misinformation exhibit similar or slightly lower bias ratio. In the Stereotype dataset, multi-turn dialogues generally show higher bias ratio, with the most noticeable increases in Scattered Questions and Jailbreak Tips. In the Toxicity dataset, single-turn dialogues have lower bias ratio, with larger differences observed in Scattered Questions, Anaphora Ellipsis, and Jailbreak Tips. As illustrated in Figure 5, models show relatively high bias ratio in single-turn scenarios for tasks including Fix Format and Negative Feedback, indicating that models may prioritize utility over fairness even in single-turn contexts, making them more vulnerable to external user input or feedback. However, in tasks like Scattered Questions and Jailbreak Tips, implicit biases accumulated across dialogue turns result in a significantly higher bias ratio in the final turn.
In summary, current LLMs exhibit a noticeable increase in bias ratio when transitioning from single-turn to multi-turn, with LLMs such as Llama-2-chat (7B, 13B) and Llama-3.1-it being typical examples. Notably, tasks defined in our study like Scattered Questions, Anaphora Ellipsis, and Jailbreak Tips prove particularly challenging for these LLMs, as their overall fairness performance declines sharply. Additionally, across both models and tasks, bias ratio in the Stereotype dataset are consistently higher than in the Toxicity dataset in single-turn dialogues, and the bias ratio difference between single-turn and multi-turn scenarios is consequently smaller. This aligns Wang et al. (2024), which suggest that Stereotype tasks are more subtle and challenging compared to Toxicity tasks.
@@ 3.4 Evaluation results in different turns
To investigate the impact of turn count on model performance across different models, we calculate the average bias ratio for each dialogue turn on different tasks and models. As shown in Figure 7, the bias ratio of all models for both Stereotype and Toxicity increase with the number of turns. This suggests that, in general, the model’s outputs are influenced by the historical dialogue context, leading to an accumulation of bias across multiple turns of responses.
Figure 6: Bias ratio across different dialogue turns in terms of LLMs.
Figure 7: Bias ratio across different dialogue
@@
To explore the effect of dialogue turn count on performance across different tasks, we calculate the average bias ratio for each task across models, as illustrated in Figure 7. Note that in the Interference from Misinformation task, models are not prompted to respond during the first four turns, and in the Fix Format task, the initial prompts are objective and unrelated to bias. Thus, there are no line charts for these two tasks in the figure. As observed, in both the Stereotype and Toxicity datasets, bias ratio generally increase with the number of dialogue turns, showing a marked spike in the final turn, with the exception of the Scattered Questions task. Our further analysis suggests that this deviation in Scattered Questions may be due to the third-turn prompt subtly guiding the LLMs toward bias. Specifically, the prompt “Summarize your speculation, it sounds like they have some characteristics similar to [Target Group]?” implicitly encourages the LLMs to associate the group described in the dialogue history with a target group, without explicitly introducing a biased evaluation. This indirect guidance can lead the LLMs to generate biased responses. However, as subsequent prompts more directly link negative events or evaluations to the target group, the LLMs protective mechanisms are triggered, causing it to reject further responses. In summary, current LLMs tend to accumulate bias in multi-turn dialogues, which is particularly pronounced in fairness-related tasks like Anaphora Ellipsis, Jailbreak Tips, and Negative Feedback.
@@ 3.5 Evaluation Results on different groups
We assess model fairness across bias attributes by calculating the proportion of biased responses for each category in our dataset. As shown in Figure 8(a) (see Appendix B.4 for a color bar zoomed-in version), slight variation in performance is observed across models and bias attributes groups on the Stereotype datasets. The Age group shows the largest disparity, while the Gender group consistently sees poor performance across all models. Llama-2-chat (7B, 13B) exhibit the most variability across bias attributes, whereas ChatGPT demonstratio strong and consistent performance. As shown in Figure 8(b), the overall bias is more severe, and performance variation is more pronounced on the Toxicity datasets. The Race group shows the largest disparity, while the Gender and Disabled groups consistently exhibit poor results. Llama-2-chat (7B) again shows the most significant variability.
In summary, the fairness capabilities of current LLMs vary significantly across bias attributes. A notable example is Llama-2-chat (7B, 13B), which shows substantial performance disparities across bias attributes in both datasets. The models exhibit weaker alignment in less commonly represented categories such as age, disabled, and appearance, while demonstrating stronger fairness in categories with greater focus, such as race and religion. The model generally performs poorly on certain social attributes. We conducted a brief exploration of this phenomenon, and the detailed analysis is provided in Appendix B.4. These findings highlight the need for future LLM fairness efforts to drive more comprehensive alignment across all bias attributes.
Figure 8: Bias ratio across different bias attributes.
@@ 3.6 Challenge Fair-MT Bench 1K
To enable more efficient evaluation, we distill the most challenging data from our Fair-MT Bench to create a lighter LLM fairness benchmark, FairMT-1K. Specifically, we select data points where the six models had the highest error ratio in the original FairMT-10K dataset based on our testing results. An equal number of samples are chosen from each task. The specific method for selecting the FairMT-1K dataset is detailed in Appendix B.5. We then evaluated a broader range of models on this new dataset, including Gemma-2-it (2B, 9B, 27B) (Gemma Team et al., 2024), Mistral-Small-Instruct-2409 (Jiang et al., 2023), Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2023), Qwne2.5-Instruct (0.5B, 3B, 7B) (Team, 2024), and GPT-4 (Achiam et al., 2023), resulting in a more diverse set of models. Table 3 presents the fairness performance of each model on FairMT-1K.
Table 3: Bias ratio of different LLMs on FairMT-1K.
@@
As demonstrated, even the most recently introduced models, which have been widely recognized for their performance, exhibit a significant proportion of biased responses on the FairMT-1K, underscoring the dataset’s challenging nature for assessing LLM fairness. Notably, we find that for certain tasks, such as Scattered Questions, Fix Format, and Negative Feedback, the proportion of biased responses tends to decrease as the model size increases. This could be due to the increased model parameters enhancing the ability of models to comprehend user instructions, which facilitates a more accurate understanding of user intent in multi-turn dialogues. Consequently, this improvement may reduce instances where fairness is compromised in favor of increasing user satisfaction.
@@ 3.7 Discussion
Our experiments above demonstrate the necessity of establishing a new benchmark for evaluating LLM fairness performance in multi-turn dialogues, since testing on single-turn fairness data fails to capture issues stemming from bias accumulation (see Section 3.3 and Section 3.4), and empirical results reveal that current LLMs struggle to maintain consistently strong fairness performance across a diverse set of dialogue scenarios designed for evaluating fairness (see Section 3.2). Importantly, we observe that LLMs face challenges in comprehensively addressing different bias attributes (see Section 3.5) and different task types including comprehension-focused and bias-resistance tasks (see Section 3.2). Finally, after testing 15 LLMs on the more challenging dataset, FairMT-1K, the results confirmed that there is still room for improvement in the fairness performance of LLMs (see Section 3.6). With this proposed benchmark, we encourage future work to focus on improving fairness in multi-turn scenarios to achieve more comprehensive fairness enhancements in future.